A gardener is putting a wire fence along the edge of his garden to keep animals from eating his plants. If he has 20 meters of fence what is the largest rectangular area he can enclose?
Answer by lwsshak3(11628)

A gardener is putting a wire fence along the edge of his garden to keep animals from eating his plants. If he has 20 meters of fence what is the largest rectangular area he can enclose?
***
let x=length of rectangular area
20-x=width of rectangular area
Area=length*width
A=x(20-x)=20x-x^2
complete the square
A=-(x^2-20x+100)+100
A=-(x-10)^2+100
This is an equation of a parabola that opens down with maximum at (10,100)
largest rectangular area gardener can enclose=100 m^2